The temperatures in Philadelphia for the past week are noted below, alongside a forecast. Which method, 2 period moving average or naive produces a better forecast based on MAD?

Actual Forecast

Today 102 100

Yesterday 99 95

Today - 2 97 95

Today - 3 94 90

Today - 4 89 85
Answer #1

Two period moving average is the average of previous period actual data

Mean absolute error

MAD = (1/n)\sum \left | Actual - Forecast \right |

Actual 2-periodErrorAbsolute error 2 Today 3 yesterday 9:9 4 Today-2 5 Today-3 6 Today-4 102 97 94 89 100.5 98 95.5 -3.5 -4 6

MAD = 4.667

Naive method:

Forecast for the present period = Actual data of the previous period.

Actual 2-period ErrorAbsolute error 2 Today 3 yesterday 9:9 4 Today-2 5 Today-3 6 Today-4 102 102 3 3 97 94 89 97 94 3 3 Tota

MAD= 3.250

naive produces a better forecast based on MAD